# Refining the Complexity Landscape of Speed Scaling: Hardness and Algorithms

### v1

@@ 1 Introduction
To formalize this trade-off, consider the following underlying scheduling problem: Given a set of $n$ preemptable jobs, each job $j$ is associated with a release time $r_{j}$ denoting the earliest time the job can begin being processed, a processing volume $v_{j}$ denoting the number of CPU cycles required for processing the job, and a weight $w_{j}$ denoting the relative importance of the job. The jobs are to be processed on a speed-scalable processor and the power-consumption corresponding to each allowable speed is given by a power function $P$. The set of different allowable speeds can be discrete or continuous (in the latter case, $P$ is usually considered to be a continuous convex function of the speed). When the processor runs at a speed $s$, it can process $s$ units of volume per unit of time. One can think of the speed as, e.g., the CPU-frequency of the processor. Naturally, the energy consumption of the processor is defined by the integral over time of $P(s(t))$, where $s(t)$ denotes the speed of the processor at time $t$. A schedule determines which job is processed at what speed at each time $t$, so that each job $j$ is fully processed after its respective release time $r_{j}$. As a QoS objective, we consider the total (weighted) flow time (also known as response time) of a schedule, that is, the (weighted) sum over all jobs of the difference between the time their processing is completed and the time they were released. In other words, the total (weighted) flow time of a schedule is the total accumulated time that jobs spend in the system. We are interested in the tradeoff between the energy consumption and the total (weighted) flow time, which we capture by considering as an objective function the sum of energy and total (weighted) flow time. [^1] Note that the desired trade-off between energy and flow can be represented by scaling the power function $P$ accordingly.
@@
Our $\mathsf{NP}$-hardness results further suggest that simultaneously optimizing the speed(s) at which each job is processed and the order at which jobs are processed is inherently difficult. For the unweighted variants, if the average speed at which each job is processed is given, then an optimal schedule can be easily constructed by prioritizing the jobs according to the shortest remaining processing time rule (SRPT). However, we show that determining the optimal speeds is $\mathsf{NP}$-hard even when given a priority ordering – for most variants of the problem even if this priority ordering corresponds to an optimal schedule. In contrast, as our second main contribution, we also show that a completion ordering is sufficient for efficiently computing optimal speed schedules. We note that such a result for FE-IDUA has been claimed before by [ComplexitySpeedScaling] as an extension of an algorithm for FE-IDUU, which was only analyzed for the case of two speeds. It is further claimed in [ComplexitySpeedScaling] for this $2$-speed algorithm that “it is straightforward to generalize it to $k$ speeds”, but we provide counterexamples for the two, in our opinion, most natural generalizations (see Section 4.1 for more details). Our algorithm on the other hand is based on a different, LP-based, approach; furthermore, our results can be extended to the budget version of each variant as well.
@@
The proof of Theorem 1.1 consists of a reduction of the budget variant of the problem B-IDUA with two available speeds, which is known to be $\mathsf{NP}$-hard [ComplexitySpeedScaling]. Via an intricate construction, we are able to simulate an energy-budget constraint within the flow-plus-energy setting for any such two-speed B-IDUA instance. At a high level, the idea is to create an additional job of comparatively large volume along with a substantial number of smaller jobs which are released significantly after all other jobs. In an optimal schedule, these latter small jobs must be processed immediately upon release so as to keep the total flow time low. This forces the the large-volume job to complete before the release of the smaller jobs, and in turn, a specific amount of volume from the original B-IDUA instance needs to be run at speed $s_{2}$ thereby simulating the original budget constraint. Our reduction gives further insight into the relationship between the corresponding budget and flow-energy trade-off variants of the problem.
@@
In addition to the results already discussed, the problem has been extended to multiprocessors (see [Bunde06], which also shows that optimal B-ICWU schedules cannot be computed by using only arithmetic operations and root computations). The online setting of different variations has also been widely explored; see [AlbersF07, BansalPS07, GuptaIKMP12, LamLTW13, BansalC09] for some examples.
@@ 2 Formal Problem Description and Notation
Consider $n$ jobs $1,\ldots,n$ that have to be scheduled on a single speed-scalable processor. Each job $j$ is associated with a corresponding release time $r_{j}\in\mathbb{R}_{\geq 0}$, a processing volume $v_{j}\in\mathbb{R}_{\geq 0}$, and weight $w_{j}\in\mathbb{R}_{\geq 0}$. The processor is equipped with $k+1$ distinct allowable nonnegative real speeds: $s_{0}<\dots<s_{k}$, with corresponding nonnegative real power consumptions $P_{0}<\dots<P_{k}$. If the processor runs at speed $s_{i}$ for $\tau$ time units, it processes a volume of $\tau\cdot s_{i}$ and consumes an amount $\tau\cdot P_{i}$ of energy. We assume that the processor idles when it uses the lowest speed (that is $s_{0}=0$), that $P_{0}=0$, and w.l.o.g., that $\min_{j}\{r_{j}\}=0$.
@@
Recall that we adopt the quintuple notation $\star$-$\star\star\star\star$ from [ComplexitySpeedScaling]. For completeness, we formally restate the relevant specific entries with the notation introduced above. For the first entry, we consider two different objectives: either minimize $F(\sigma)+E(\sigma)$, or minimize only $F(\sigma)$ subject to an energy constraint $E(\sigma)\leq B$ for some given budget $B$. We refer to the former as the flow + energy variant (FE), and the latter as the budget variant (B). For the second entry, we only consider integral flow (I) $F(\sigma)$ as defined above, but the related concept of a fractional flow (F) has been considered in the literature. For the third entry, the allowed speeds of the processor could either be discrete (D) as defined above, or continuous (C), where the set of allowable speeds is $[0,H)$ for some $H>0$, or $0,\infty)$ and the power consumption is defined by a function $P$ of the speed with $P(0)=0$. Note that Observation 2.2 implies that the discrete speed setting can be modeled by the continuous one with a piecewise linear power function $P$ and the discrete speed setting is therefore a special case of the continuous setting (see also Theorem 7 in [ComplexitySpeedScaling]). For the fourth entry, in the unweighted variant (U), we have $w_{1}=\dots=w_{n}=1$, whereas for the weighted variant (W) we only require that $w_{j}\in\mathbb{R}_{>0}$ for all $j\in\{1,\dots,n\}$. Finally, for the fifth entry, in the unit size variant (U), we have $v_{1}=\dots=v_{n}=1$, whereas in the arbitrary size variant (A) we only require that $v_{j}\in\mathbb{R}_{>0}$ for all $j\in\{1,\dots,n\}$. Recall that we also consider each variant with a fixed priority (P) or completion (C) ordering and extend the quintuple notation by a suffix accordingly.
@@ 2.1 Affection and Shrinking Energy
In this section we introduce the notions of affection and shrinking energy. Intuitively, the affection of a job measures the improvement in total flow resulting from processing that job in less amount of time, while its shrinkage energy measures the energy needed to achieve this reduction in the processing time of the job. The affection and shrinking energy combined, provide information on which jobs (if any) one could accelerate in order to obtain an improved schedule. Similar definitions are used in [ComplexitySpeedScaling], however, we do deviate from their notation, so as to be able to use it in a more general setting.
@@ 3.1 Hardness of FE-IDUA
Figure 3: If we take an optimal schedule for $\mathcal{I}^{B}$ and add the jobs $n+1,\dots,2n+1$ at speed $s_{1}$, then we obtain a schedule for $\mathcal{I}^{FE}$ that satisfies the following. Jobs $1,\dots,n$ have a total processing time of VYV-Y, and are the only jobs that can run at a speed higher than $s_{1}$.
@@ 3.2 Hardness of FE-IDWU
Since $A<ma_{1}\leq m^{2}a_{1}^{2}$ we have that $K=\left\lceil\frac{m^{2}}{2}+\frac{A}{2a_{1}^{2}}\right\rceil\leq\left\lceil\frac{m^{2}}{2}+\frac{m^{2}}{2}\right\rceil=m^{2}\penalty 10000\$. W.l.o.g. we assume that $m\geq 2$. Using that $r_{m+1}=m(m+1)+K-Y$, we obtain
@@ 3.3 Fixed Priority Ordering
Consider an instance $\mathcal{I}^{B}$of B-IDUA-P with priority ordering $1\prec\dots\prec n$, and assume this ordering is known to be optimal. Let $\mathcal{I}^{FE}$ be the FE-IDUA instance obtained by applying the reduction from Section 3.1 to $\mathcal{I}^{B}$. Let $n+1,\dots,2n+1$ be the newly added jobs, as described in the reduction. By Lemma 3.1 and the proof of Lemma 3.5, any optimal schedule for $\mathcal{I}^{FE}$adheres to the priority ordering $1\prec\dots\prec n\prec n+1\prec\dots\prec 2n+1\penalty 10000\$. From Lemma 3.7 and Theorem 3.20, it follows that FE-IDUA-P is $\mathsf{NP}$-hard, even when the priority ordering is known to be optimal.
@@ 4 An LP for Fixed Completion Ordering
  • $\hat{C}_{j}(\sigma):=\max\{C_{j}(\sigma),\hat{C}_{j-1}(\sigma)\}$$\forall j\in\{2,\dots,n\},$
@@
4.1 Towards a Combinatorial Algorithm for FE-ID**-C
@@
Here we provide a different and simple combinatorial algorithm for FE-IDUU with an arbitrary number of speeds. In fact, Algorithm 1 is a greedy algorithm that increases the speed of the job for which that increase in speed most strongly improves the objective function. Furthermore, we conjecture a slight adaptation of this algorithm also solves any *-ID**-C variant.
@@
Algorithm 1 computes an optimal schedule for FE-IDUU.
Proof 4.8.
@@
Recall the definition of affection and note that Observation 2.4 does not hold for $\hat{F}(\sigma)=\sum_{j}{\hat{C}_{j}(\sigma)-r_{j}}$ instead of $F(\sigma)$. For a fixed completion ordering we define the extended affection as follows.
@@
  • The extended affection of job $j$ is defined as $\hat{\kappa}_{j}(\sigma):=\sum_{j^{\prime}\in\hat{K}_{j}(\sigma)}w_{j^{\prime}}\penalty 10000\$.
@@
The algorithm obtained by replacing $\kappa$ with $\hat{\kappa}$ in Algorithm 1 computes an optimal schedule for any FE-ID**-C variant.
@@ 5 Discussion and Future Work
We showed $\mathsf{NP}$-hardness of FE-IDUA, FE-ICUA, FE-IDWU and FE-ICWU, and furthermore that all these problem variants are solvable in polynomial time when given a completion-time ordering.
@@ Appendix B Hardness of B-IDUA-P
$\leq(1+4\delta)\sum_{i=1}^{m}a_{i}+\sum_{i\notin L}a_{i}-2\delta a_{i}-\varepsilon\leq B-\sum_{i\in L}2\delta a_{i}-\varepsilon$
@@ C.1 Unsuccessful FE-IDUU-Algorithm Generalizations to 3+ Speeds
Consider an instance of FE-IDUU with two speeds: $s_{1}$ and $s_{2}$. For jobs $1,\dots,n$, we assume that $r_{1}\leq\dots\leq r_{n}$. Since we have unit size and unit weight jobs, ordering the jobs from earliest to latest release time (FIFO) gives an optimal completion ordering. In [ComplexitySpeedScaling] it is shown that the following algorithm computes optimal schedules for this variant. For ease of notation, we use $\kappa_{j}$ and $\Delta_{j}$ to denote the affection and shrinking energy of $j$ during the algorithm (even though these values might change throughout the algorithm).
@@
Note that once a job $j$ reaches speed $s_{2}$, we have $\Delta_{j}=\infty$, and thus the algorithm will not speed $j$ up any further.
Now consider an instance of FE-IDUU with an arbitrary number $k$ of speeds, say $s_{1},\dots,s_{k}$. In our opinion the following are the most natural generalizations of the two-speed algorithm. In the first one, simply run Algorithm 2. This means even if when a job $j$ fully runs at $s_{2}$, and $\Delta_{j}$ has increased to $\Delta_{2}$, the algorithm would still continue increasing the speed of $j$ if it still holds that $\kappa_{j}\geq\Delta_{j}$.

### v2

@@ 1 Introduction
To formalize this trade-off, consider the following underlying scheduling problem: Given a set of $n$ preemptable jobs, each job $j$ is associated with a release time $r_{j}$ denoting the earliest time the job can begin being processed, a processing volume $v_{j}$ denoting the number of CPU cycles required for processing the job, and a weight $w_{j}$ denoting the relative importance of the job. The jobs are to be processed on a speed-scalable processor and the power-consumption corresponding to each allowable speed is given by a power function $P$. The set of different allowable speeds can be discrete or continuous (in the latter case, $P$ is usually considered to be a continuous convex function of the speed). When the processor runs at a speed $s$, it can process $s$ units of volume per unit of time. One can think of the speed as the CPU frequency. Naturally, the energy consumption of the processor is defined by the integral over time of $P(s(t))$, where $s(t)$ denotes the speed of the processor at time $t$. A schedule determines which job is processed at what speed at each time $t$, so that each job $j$ is fully processed after its respective release time $r_{j}$. As a QoS objective, we consider the total (weighted) flow time (also known as response time) of a schedule, that is, the (weighted) sum over all jobs of the difference between the time their processing is completed and the time they were released. In other words, the total (weighted) flow time of a schedule is the total accumulated time that jobs spend in the system. We are interested in the tradeoff between the energy consumption and the total (weighted) flow time, which we capture by considering as an objective function the sum of energy and total (weighted) flow time. [^1] Note that the desired trade-off between energy and flow can be represented by scaling the power function $P$ accordingly.
@@
Our $\mathsf{NP}$-hardness results further suggest that simultaneously optimizing the speed(s) at which each job is processed and the order at which jobs are processed is inherently difficult. For the unweighted variants, if the average speed at which each job is processed is given, then an optimal schedule is easily constructed by prioritizing the jobs according to the shortest remaining processing time rule (SRPT). However, we show that determining the optimal speeds is $\mathsf{NP}$-hard even when given a priority ordering – for most variants of the problem even if this priority ordering corresponds to an optimal schedule. In contrast, as our second main contribution, we also show that a completion ordering is sufficient for efficiently computing optimal speed schedules. We note that such a result for FE-IDUA was claimed before by Barcelo et al. [ComplexitySpeedScaling] as an extension of an algorithm for FE-IDUU, which was only analyzed for the case of two speeds. They claim that, for this $2$-speed algorithm, “it is straightforward to generalize it to $k$ speeds”, but we provide counterexamples for the two, in our opinion, most natural generalizations (see Section 5 for more details). Our algorithm on the other hand uses a different, LP-based approach. Furthermore, our results can be extended to the budget version of each variant as well.
@@
The proof of Theorem 1.1 consists of a reduction from the budget variant of the problem B-IDUA with two available speeds, which is known to be $\mathsf{NP}$-hard [ComplexitySpeedScaling]. Via an intricate construction, we are able to simulate an energy-budget constraint within the flow-plus-energy setting for any such two-speed B-IDUA instance. At a high level, the idea is to create an additional job of comparatively large volume along with a substantial number of smaller jobs which are released significantly after all other jobs. In an optimal schedule, these latter small jobs must be processed immediately upon release so as to keep the total flow time low. This forces the large-volume job to complete before the release of the smaller jobs, and in turn, a specific amount of volume from the original B-IDUA instance needs to be run at speed $s_{2}$ thereby simulating the original budget constraint. Our reduction gives further insight into the relationship between the corresponding budget and flow-energy trade-off variants of the problem.
@@
In addition to the results already discussed, the problem has been extended to multiprocessors (see [Bunde06], which also shows that exact optimal B-ICWU schedules cannot be computed even on a machine that can do real arithmetic operations and root computations). The online setting of different variations has also been widely explored; see [AlbersF07, BansalPS07, GuptaIKMP12, LamLTW13, BansalC09] for some examples.
@@ 2 Formal Problem Description and Notation
Consider $n$ jobs $1,\ldots,n$ that have to be scheduled on a single speed-scalable processor. Each job $j$ is associated with a corresponding release time $r_{j}\in\mathbb{Q}_{\geq 0}$, a processing volume $v_{j}\in\mathbb{Q}_{\geq 0}$, and weight $w_{j}\in\mathbb{Q}_{\geq 0}$. The processor is equipped with $k+1$ distinct nonnegative real speeds: $s_{0}<\dots<s_{k}$, with corresponding nonnegative real power consumptions $P_{0}<\dots<P_{k}$. If the processor runs at speed $s_{i}$ for $\tau$ time units, it processes a volume of $\tau\cdot s_{i}$ and consumes an amount $\tau\cdot P_{i}$ of energy. We assume that the processor idles when it uses the lowest speed (that is $s_{0}=0$), that $P_{0}=0$, and w.l.o.g., that $\min_{j}\{r_{j}\}=0$.
@@
Recall that we adopt the quintuple notation $\star$-$\star\star\star\star$ from [ComplexitySpeedScaling]. For completeness, we formally restate the relevant specific entries with the notation introduced above. For the first entry, we consider two different objectives: either minimize $F(\sigma)+E(\sigma)$, or minimize only $F(\sigma)$ subject to an energy constraint $E(\sigma)\leq B$ for some given budget $B$. We refer to the former as the flow + energy variant (FE), and the latter as the budget variant (B). For the second entry, we only consider integral flow (I) $F(\sigma)$ as defined above, but the related concept of a fractional flow (F) has been considered in the literature. For the third entry, the allowed speeds of the processor could either be discrete (D) as defined above, or continuous (C), where the set of allowable speeds is $[0,H)$ for some $H>0$, or $0,\infty)$ and the power consumption is defined by a function $P$ of the speed with $P(0)=0$. Note that Observation 2.2 implies that the discrete speed setting can be modeled by the continuous one with a piecewise linear power function $P$ and the discrete speed setting is therefore a special case of the continuous setting (see also Theorem 7 in [ComplexitySpeedScaling]). For the fourth entry, in the unweighted variant (U), we have $w_{1}=\dots=w_{n}=1$, whereas for the weighted variant (W) we only require that $w_{j}\in\mathbb{Q}_{>0}$ for all $j\in\{1,\dots,n\}$. Finally, for the fifth entry, in the unit size variant (U), we have $v_{1}=\dots=v_{n}=1$, whereas in the arbitrary size variant (A) we only require that $v_{j}\in\mathbb{Q}_{>0}$ for all $j\in\{1,\dots,n\}$. Recall that we also consider each variant with a fixed priority (P) or completion (C) ordering and extend the quintuple notation by a suffix accordingly.
@@ 2.1 Affection and Shrinking Energy
In this section we introduce the notions of affection and shrinking energy. Intuitively, the affection of a job measures the improvement in total flow resulting from processing that job in less time (i.e., at higher speed), while its shrinkage energy measures the energy needed to achieve this reduction in the processing time of the job. The affection and shrinking energy combined, provide information on which jobs (if any) one could accelerate in order to obtain an improved schedule. Similar definitions are used in [ComplexitySpeedScaling], however, we do deviate from their notation, so as to be able to use it in a more general setting.
@@ 3.1 Hardness of FE-IDUA
Figure 3: Jobs $n+1,\dots,2n+1$ are added to an optimal schedule for $\mathcal{I}^{B}$ and run at speed $s_{1}$. This gives a schedule for $\mathcal{I}^{FE}$ where jobs $1,\dots,n$ have a total processing time of VYV-Y, and are the only jobs that run at a speed higher than $s_{1}$.
@@ 3.2 Hardness of FE-IDWU
Since $A<ma_{1}\leq m^{2}a_{1}^{2}$ we have that $K=\left\lceil\frac{m^{2}}{2}+\frac{A}{2a_{1}^{2}}\right\rceil\leq\left\lceil\frac{m^{2}}{2}+\frac{m^{2}}{2}\right\rceil=m^{2}$. W.l.o.g., we assume that $m\geq 2$. Using that $r_{m+1}=m(m+1)+K-Y$, we obtain
@@ 3.3 Fixed Priority Ordering
Consider an instance $\mathcal{I}^{B}$of B-IDUA-P with priority ordering $1\prec\dots\prec n$, and assume this ordering is known to be optimal. Let $\mathcal{I}^{FE}$ be the FE-IDUA instance obtained by applying the reduction from Section 3.1 to $\mathcal{I}^{B}$. Let $n+1,\dots,2n+1$ be the newly added jobs, as described in the reduction. By Lemma 3.1 and the proof of Lemma 3.5, any optimal schedule for $\mathcal{I}^{FE}$adheres to the priority ordering $1\prec\dots\prec n\prec n+1\prec\dots\prec 2n+1$. From Lemma 3.7 and Theorem 3.20, it follows that FE-IDUA-P is $\mathsf{NP}$-hard, even when the priority ordering is known to be optimal.
@@ 4 An LP for Fixed Completion Ordering
$\hat{C}_{j}(\sigma)$ $:=\max\{C_{j}(\sigma),\hat{C}_{j-1}(\sigma)\}\,,\quad\forall j\in\{2,\dots,n\}\,$,
@@
5 Towards a Combinatorial Algorithm for FE-ID**-C
@@
Here, we provide a different and simple combinatorial algorithm for FE-IDUU with an arbitrary number of speeds. For unit-size and unit-weight jobs, ordering the jobs from earliest to latest release date (FIFO) gives an optimal completion ordering for any speed profile $S(t)$. This follows from a simple exchange argument. Hence, the problem is equivalent to finding optimal speeds. Algorithm 1 is a greedy algorithm that increases the speed of a job that leads to the biggest improvement in the objective function. Furthermore, we conjecture a slight adaptation of this algorithm also solves any *-ID**-C variant.
@@
   Increase the speed of job $j^{*}:=\arg\max_{j}(\kappa_{j}-\Delta_{j})$ (break ties arbitrarily), until either $\kappa_{j^{*}}-\Delta_{j^{*}}<0$, or $\kappa_{j^{*}}-\Delta_{j^{*}}\neq\max_{j}(\kappa_{j}-\Delta_{j})$.
@@
Algorithm 1 computes an optimal schedule for FE-IDUU in polynomial time.
Proof 5.2.
@@
Recall the definition of affection and note that Observation 2.4 does not hold for $\hat{F}(\sigma)=\sum_{j}(\hat{C}_{j}(\sigma)-r_{j})$ instead of $F(\sigma)$. For a fixed completion ordering we define the extended affection as follows.
@@
  • The extended affection of job $j$ is defined as $\hat{\kappa}_{j}(\sigma):=\sum_{j^{\prime}\in\hat{K}_{j}(\sigma)}w_{j^{\prime}}$.
@@
The algorithm obtained by replacing $\kappa$ with $\hat{\kappa}$ in Algorithm 1 computes an optimal schedule for any FE-ID**-C variant in polynomial time.
@@ 6 Discussion and Future Work
We showed the $\mathsf{NP}$-hardness of FE-IDUA, FE-ICUA, FE-IDWU and FE-ICWU, and furthermore that all these problem variants are solvable in polynomial time when given a completion-time ordering.
@@ Appendix B Hardness of B-IDUA-P
$\leq(1+4\delta)\sum_{i=1}^{m}a_{i}+\sum_{i\notin L}a_{i}-2\delta a_{i}-\varepsilon\leq B-\sum_{i\in L}2\delta a_{i}-\varepsilon\,.$
@@ Appendix C Towards a Combinatorial Algorithm for FE-ID**-C
Consider an instance of FE-IDUU with jobs $1,\dots,n$ and speeds $s_{1},\dots,s_{k}$. We assume that $r_{1}\leq\dots\leq r_{n}$. Since we have unit-size and unit-weight jobs, ordering the jobs from earliest to latest release date (FIFO) gives an optimal completion ordering for any speed profile $S(t)$. This follows from a simple exchange argument. Hence, the problem is equivalent to finding optimal speeds for completion ordering $1\prec\dots\prec n$. In this section, we use only job speeds to describe a schedule.
@@ C.1 Unsuccessful FE-IDUU-Algorithm Generalizations to 3+ Speeds
Barcelo at al. [ComplexitySpeedScaling] show that the following algorithm computes optimal schedules for this variant. For ease of notation, we use $\kappa_{j}$ and $\Delta_{j}$ to denote the affection and shrinking energy of $j$ during the algorithm (even though these values change throughout the algorithm).
@@
Note that, once a job $j$ reaches speed $s_{2}$, we have $\Delta_{j}=\infty$, and thus the algorithm will not speed $j$ up any further.
Now consider an instance of FE-IDUU with an arbitrary number $k$ of speeds, say $s_{1},\dots,s_{k}$. In our opinion the following are the most natural generalizations of the two-speed algorithm. In the first one, simply run Algorithm 2. This means that, even if a job $j$ fully runs at $s_{2}$, and $\Delta_{j}$ has increased to $\Delta_{2}$, the algorithm increases the speed of $j$ as long as $\kappa_{j}\geq\Delta_{j}$ holds.